Consider explaining the selling prices of condos (price) using unit size (sq_feet), number of bathrooms (num_bath), quality of finishes index (fin_index), level (level), and distance to city centre (dist_centre_km). For observational, cross-sectional data with 66 observations, what is the critical value for the overall test of statistical significance at a 10% significance level? (Record your answer accurate to at least the nearest second decimal place with standard rounding.)
Answer #1

here degree of freedom for numerator p=number of independent variables =5

degree of freedom for Denominator=n-p-1=66-5-1 =60

therefore

critical value at 10% level and (5,60) degree of freedom =1.95